# Public Projects with Preferences and Predictions

### v1

@@ 3.2 Analysis
Budget balance follows immediately from budget balance of the QTM.
For Price of Anarchy, we first analyze a hypothetical execution of the original QTM with synthetic and real players. Then, we use properties of that game to analyze the Impractical Variant of the Synthetic Players QTM (“our mechanism”).
Given $\mathbf{v}$, let $c\geq\tfrac{1}{2}\max_{i,k}v_{k}^{i}$. Create a set of synthetic players as follows: choose $\hat{n}\geq\frac{\max_{k}B_{k}}{2c}$ and for $j=1,\dots,\hat{n}$, define $\mathbf{\hat{v}}^{j}=\tfrac{1}{\hat{n}}\mathbf{B}$. Observe $c\geq\tfrac{1}{2}\max_{j,k}\hat{v}_{k}^{j}$.
[…]
Now, return to our mechanism. An agent $i$’s best-response depends only on $\mathbf{v}^{i}$ and on $\mathbf{a}^{\text{mech}}+\sum_{j\neq i}\mathbf{a}^{j}$, the vector of total votes excluding $i$’s. This follows because the terms of $i$’s utility that $i$ controls depends only on these quantities. So for all real agents, the utility and best responses are strategically equivalent in our mechanism and in the synthetic game where the synthetic players commit to total votes $\mathbf{a}^{\text{mech}}$. We have already shown that at least one such equilibrium $\mathbf{a}$ exists, with the vote totals $\mathbf{A}$ computed and announced by the mechanism. By the focal equilibrium assumption, the real agents play that equilibrium $\mathbf{a}$, in which case the output $\mathbf{p}$ is identical to that of the synthetic game on equilibrium $(\mathbf{a},\mathbf{a}^{\text{mech}})$. Furthermore, in every strategy profile $\mathbf{a}$ of our mechanism, social welfare $\sum_{k}p_{k}W_{k}$ is equal to the social welfare of the synthetic game in strategy profile $(\mathbf{a},\mathbf{a}^{\text{mech}})$. By Theorem 1, in any pure-strategy equilibrium of the synthetic game, including those where the synthetic players play $\mathbf{a}^{\text{mech}}$, welfare is at least $W_{1}\cdot\max\{\tfrac{1}{2},1-(2/T)^{2/5}\}$. ∎
4 Incorporating Predictions
We now consider public projects with predictions. Specifically, we propose that an information-aggregation mechanism (just called an aggregation mechanism) be operated first in order to advise the decisionmaking mechanism, which operates second. We refer to participants in the first stage as aggregators and participants in the second as decisionmakers.
Incentives for manipulation.
If the aggregators and decisionmakers are disjoint, then we can simply use any aggregation mechanism to obtain estimates of the external welfare impact $B_{k}$ of each project $k$, then use the Synthetic Players QTM previously proposed. The challenge comes when agents may participate in both mechanisms. There are two potential problems: a decisionmaker may manipulate the predictions in order to boost some alternatives’ chances of being chosen, and an aggregator may manipulate the decisionmaking mechanism in order to increase the rewards for their predictions.
4.1 General Mechanism
Formally, we model an aggregation mechanism as a game that produces predictions of the external welfare impacts of each alternative. We let $\hat{B}_{k}$ denote the produced estimate of the external welfare impact of alternative $k$, should it be chosen. The aggregation mechanisms we consider, prediction markets and wagering mechanisms, reward participants based on an observation of the actual outcome once an alternative is chosen. The combined mechanism we propose, the Synthetic players QUAdratic transfer mechanism with Predictions (SQUAP), is as follows:
  1. 1.
    Agents participate in the aggregation mechanism, which produces $\mathbf{\hat{B}}=(\hat{B}_{1},\dots,\hat{B}_{m})$.
  2. 2.
    We run the Synthetic Players QTM defined in Section 3.1 using external welfare impacts $\mathbf{\hat{B}}$. As discussed below, instead of redistributing payments, we may use them to fund the aggregation mechanism.
  3. 3.
    The aggregation mechanism observes the choice of alternative and subsequent events and rewards its participants.
Again, we do not know how to analyze SQUAP, and we turn to Impractical SQUAP, which is the same except:
  • Instead of using the Synthetic Players QTM in the decisionmaking stage, we use the Impractical Variant of the Synthetic Players QTM, which assumes knowledge of the vote totals.
  • We do not redistribute payments from the QTM.
Redistribution has the potential at least in theory to cause potential incentive problems (see Lemma 3). For simplicity, we generally assume that the mechanism keeps or “burns” the payments, or uses them to subsidize the aggregation mechanism. In practice, designers may choose to go ahead with the redistribution, as the incentive concern is likely low.
Incentive alignment.
We emphasize that any number of agents may participate in both mechanisms or in either one alone. Perhaps surprisingly, we only require two natural properties of an aggregation mechanism in order to prove welfare guarantees in equilibrium of the combined mechanism. These properties address the two possible incentive problems discussed above. First is the incentive of aggregators to manipulate the decision in order to gain higher prediction rewards.
Definition 1.
An aggregation mechanism is alternative-independent if a participant’s net expected payment does not depend on which alternative $k$ is selected by the decisionmaking mechanism.
We will see that an importance-weighting technique due to Chen et al. [2011] allows one to achieve alternative-independence, so that prediction rewards are independent of the decisionmaking mechanism (at least in expectation).
Next is the incentive of decisionmakers to manipulate the predictions $\mathbf{\hat{B}}$ in order to boost the chances of their preferred alternative(s). In general, agents can gain from manipulating the estimated external welfare even at a cost. However, we will be able to show that the decisionmaking stage’s welfare guarantees are robust to relatively small manipulations. Therefore, we only need that large manipulations are prohibitively costly. We formalize this costliness as follows. Say that a player in a game is $x$-best-responding if they can improve their net expected payoff by at most $x$ by switching to another strategy.
Definition 2.
In an aggregation mechanism, let $\mathbf{B}$ be the output of the mechanism if all agents are truthful and let $\mathbf{\hat{B}}$ be the output under strategic behavior. Given some $x>0$ and $\alpha>0$, we say the aggregation mechanism has a deviation bound $\alpha x$ if, for all strategy profiles where all participants are $x$-best-responding,
$\max_{k}|\hat{B}_{k}-B_{k}|\leq\alpha x.$
To quantify the welfare of the combined mechanism, we assume that $\mathbf{B}$, the output of the aggregation mechanism if all agents are truthful, is the true external welfare impact. Therefore, the total welfare of alternative $k$ is $W_{k}=B_{k}+V_{k}$, and we again assume $W_{1}\geq\cdots\geq W_{m}$. Because this is now a dynamic game, we also consider a small refinement of equilibrium: We say the combined mechanism is in sub-mechanism perfect pure equilibrium if it is in Nash equilibrium and, in addition, the strategies played at each subgame consisting of the Synthetic Players QTM are in pure-strategy Nash equilibrium. We also continue to assume agents play the focal pure-strategy equilibrium in the Impractical Synthetic Players QTM. In an abuse of notation, we let $\text{pPoA}(\mathbf{v},\mathbf{B})$ refer to the ratio of the welfare in the worst such sub-mechanism perfect pure equilibrium to $W_{1}$.
@@ 4.1 General Mechanism
Define the spread of the game to be $T\coloneqq\frac{W_{1}}{max_{i,k}v_{k}^{i}}$. Suppose we run Impractical SQUAP with an aggregation mechanism that is alternative-independent and has a deviation bound $\alpha\max_{i,k}v_{k}^{i}$. Then in the two-alternative setting, there exists a choice of $c$ such that
[…]
The proof has two main steps. First, we use our strategic properties of alternative-independence and the deviation bound to show that, in equilibrium, the estimates $\mathbf{\hat{B}}$ are reasonably accurate. (Actually, their accuracy can even degrade somewhat as the spread of the game increases and pPoA will still tend to one.) Then, we modify our QTM Price of Anarchy analysis to obtain the welfare guarantee.
@@
In any sub-mechanism perfect pure equilibrium of SQUAP, if the aggregation mechanism is alternative-independent and has a deviation bound $\alpha x$, then its output satisfies for all $k$ the bound $|\hat{B}_{k}-B_{k}|\leq\alpha\max_{i,k}v_{k}^{i}$.
Proof.
Fix any events of the aggregation mechanism and consider the subgame of the QTM. Let an agent’s net utility in the combined mechanism be $f(k)=g(k)+h(k)$ if alternative $k$ is selected, where $g(k)$ is the net payoff from the aggregation mechanism and $h(k)$ is the net utility in the QTM stage, i.e. $v_{k}^{i}-c\sum_{\ell}(a_{\ell}^{i})^{2}$. By alternative-independence, $\operatorname{\mathbb{E}}[g(k)]=C$ for some constant $C$ independent of $k$. By design of the QTM without redistribution, $\max_{k}h(k)\leq\max_{i,k}v_{k}^{i}$. It immediately follows that in any QTM subgame equilibrium, all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding.
Now we aim to show that that in any combined equilibrium, in the aggregation stage all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding. We proceed by contradiction: assume not. Then there exists an agent in this equilibrium who is $x$-best-responding in the aggregation mechanism for $x>\max_{i,k}v_{k}^{i}$. By definition, the agent can deviate to a strategy in the aggregation mechanism which leads to a utility increase strictly larger than $\max_{i,k}v_{k}^{i}$. Moreover, if the agent deviates to casting zero votes in the QTM, she loses at most $\max_{i,k}v_{k}^{i}$ in utility, since $f(k)-f(k^{\prime})\leq\max_{i,k}v_{k}^{i}$ for any pair of alternatives $k,k^{\prime}$. It follows that this deviation strategy leads to a strictly positive change in utility; thus, the current strategy is not an equilibrium. By contradiction, then, the statement holds.
Since all agents are $(\max_{i,k}v_{k}^{i})$-best responding, the deviation bound implies the result. ∎
We now prove Theorem 2, the Price of Anarchy bound. The general strategy mirrors our QTM analysis, where we first show that $p_{1}$ approaches $1$ for a large gap $W_{1}-W_{2}$.
Lemma 4.
If the aggregation mechanism is alternative-independent and has deviation bound $\alpha x$, then in the two alternative setting, if $W_{1}-W_{2}\geq 2\alpha\max_{i,k}v_{k}^{i}$, then in any submechanism-perfect pure equilibrium,
@@
Define $\hat{W}_{k}=V_{k}+\hat{B}_{k}$, and recall $W_{k}=V_{k}+B_{k}$. By Lemma 3, we have
$\hat{W}_{1}-\hat{W}_{2}$ $=W_{1}-W_{2}+\hat{B}_{1}-B_{1}+\hat{B}_{2}-B_{2}$
@@
In the QTM stage, the synthetic players’ aggregate values are $\mathbf{\hat{B}}$. Therefore, by Proposition 4, it satisfies
@@
$=1-\left(\frac{8c}{Y}\right)^{2/3}$
@@ 4.2 Prediction Markets
First, we consider using prediction markets to aggregate information on external welfare in the first stage. Prediction markets are financial markets designed specifically for aggregating the predictions of agents by rewarding accuracy once an outcome is observed. We briefly define scoring-rule based prediction and decision markets below, referring the reader to references in Section 1.2 for more background. For simplicity, we will suppose that agents predict $\mathbf{B}$ directly and the mechanism is later able to exactly observe $B_{k}^{*}$, the true external welfare of the alternative $k$ that was selected. One could consider markets that are more complex and predict other relevant variables, either because observing $B_{k}^{*}$ directly is difficult, or to facilitate information aggregation. In this case, the mechanism could compute an estimate of $\mathbf{B}$ from the prediction markets’ forecasts and use the estimate in SQUAP.
@@
We define a scoring-rule prediction market for a real-valued random variable, following e.g. Abernethy et al. [2013]. The market defines an initial estimate $\hat{b}^{0}$. Agents arrive in sequence and provide estimates $\hat{b}^{1},\dots,\hat{b}^{N}$. Later, when the outcome $b^{*}$ of the random variable is observed, the provider of each estimate $\hat{b}^{t}$ receives a net payoff $s(\hat{b}^{t},b^{*})-s(\hat{b}^{t-1},b^{*})$, where $s$ is some proper scoring rule for the mean [Savage, 1971, Gneiting and Raftery, 2007]: the expected value of $s(\hat{b},b^{*})$, over draws $b^{*}$ from some distribution $D^{*}$, is maximized by choosing $\hat{b}=\operatorname{\mathbb{E}}[b^{*}]$. In particular, here we we will focus on the quadratic score, with a parameter $\beta>0$:
$s(\hat{b},b^{*})=-\frac{1}{\beta}(\hat{b}-b^{*})^{2}.$ (9)
Observe that each participant’s net payoff can be positive or negative, and the total payment of the mechanism telescopes to $s(\hat{b}^{N},b^{*})-s(\hat{b}^{0},b^{*})$. Following the prediction market literature, we refer to $\beta$ as the liquidity parameter. $\beta$ determines how sensitive the payoffs are to small changes in a prediction. If $\beta$ is small, then the total payoffs available are very large, and a significant payment may be required for changing the prediction even a small amount.
@@
In a decision market (e.g. [Chen et al., 2011]), we simultaneously run $m$ prediction markets, one for each of $m$ alternatives. Here, we model this as each participant $t=1,\dots$ providing an estimate $\mathbf{\hat{b}}^{t}\in\operatorname{\mathbb{R}}^{m}$. Then, we make a decision $k$ based on the outcome of the markets, and later observe $b_{k}^{*}$. We cancel all trades in the other $m-1$ markets, i.e. we only assign payoffs based on predictions $\hat{b}_{k}^{t}$ made in the $k$th prediction market.
In general, decision markets can have incentives to misreport and manipulate the decision. If the decision $k$ is made from a full-support probability distribution $\mathbf{p}$, as in the QTM, then we can address this issue with the importance-weighted or inverse-weighted scoring rule proposed by Chen et al. [2011], specialized to the quadratic score:
$S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k}^{*})=\frac{1}{p_{k}}s(\hat{b}^{t}_{k},b_{k}^{*}).$
This rule can be impractical when probabilities are small, and practitioners may prefer to replace with it something bounded, e.g. if ex post budget balance is required. However, the theoretical benefit is that, even when $\mathbf{p}$ is chosen by some rule that depends arbitrarily on the predictions, there is no incentive to manipulate predictions to increase expected score. Specifically, if the outcome $b_{k}^{*}$ is drawn from a distribution $D_{k}^{*}$, then the expected score over both $k\sim\mathbf{p}$ and $b_{k}^{*}\sim D_{k}^{*}$ is
$\sum_{k}p_{k}\operatorname{\mathbb{E}}_{D_{k}^{*}}S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k}^{*})$ $=\frac{-1}{\beta}\sum_{k}\operatorname{\mathbb{E}}_{D_{k}^{*}}(\hat{b}^{t}_{k}-b_{k}^{*})^{2}$
$=\frac{-1}{\beta}\sum_{k}\left((\hat{b}^{t}_{k}-B_{k})^{2}+\text{Var}(D_{k}^{*})\right),$
where $B_{k}=\operatorname{\mathbb{E}}_{D_{k}^{*}}[b_{k}^{*}]$. The variance term is independent of any agents’ actions, and moreover cancels out in an agent’s payoff when we consider the market scoring rule (difference of two scores). Therefore, under the importance-weighted scoring rule, it is without loss of generality to define the expected score of a prediction $\mathbf{\hat{b}}$, given the vector of means $\mathbf{B}=B_{1},\dots,B_{m}$, to be
$S(\mathbf{\hat{b}},\mathbf{B})$ $=\frac{-1}{\beta}\sum_{k}(\hat{b}_{k}-B_{k})^{2}.$ (10)
Observation 1.

### v2

@@ 5.2 SQUAP Price of Anarchy
Fix any events of the aggregation mechanism and consider the subgame of the QTM. Recall that an agent’s expected utility in the combined mechanism is $u^{i}(\mathbf{\hat{b}},\mathbf{a},\mathbf{p})=u^{i,1}(\mathbf{\hat{b}}^{i})+u^{i,2}(\mathbf{a},\mathbf{p})$. By alternative-independence, $u^{i,1}(\mathbf{\hat{b}}^{i})=C$ for some constant $C$ independent of $\mathbf{p}$. Meanwhile, by design of the QTM without redistribution, $u^{i,2}(\mathbf{a},\mathbf{p})\leq\max_{i,k}v_{k}^{i}$ for any $\mathbf{p}$. It immediately follows that in any QTM subgame equilibrium, all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding.
Now we aim to show that in any combined equilibrium, in the aggregation stage all agents are ($\max_{i,k}v_{k}^{i}$)-best-responding. We proceed by contradiction: assume there is some agent who is not. By definition, the agent can deviate to a strategy in the aggregation mechanism which leads to a utility increase strictly larger than $\max_{i,k}v_{k}^{i}$. Moreover, if the agent deviates to voting $\hat{\mathbf{a}}^{i}=\mathbf{0}$ in the QTM, she loses at most $\max_{i,k}v_{k}^{i}$ in utility because (1) $u^{i,1}$ cannot change by alternative independence and (2) $u^{i,2}(\mathbf{a},\mathbf{p})-u^{i,2}(\mathbf{a}^{-i},\hat{\mathbf{a}}^{i},\mathbf{\hat{p}})\leq\max_{i,k}v_{k}^{i}$ for any two vectors $\mathbf{p},\mathbf{\hat{p}}$. It follows that this deviation strategy leads to a strictly positive change in utility; thus, the current strategy is not an equilibrium. By contradiction, then, the statement holds. Since all agents are $(\max_{i,k}v_{k}^{i})$-best responding and the aggregation mechanism is $(\alpha,x)$-robust, the lemma follows. ∎
We are now able to bound on the Price of Anarchy in the $m=2$ case by modifying our QTM Price of Anarchy analysis. Note that the incentive for manipulation, captured by the size of $\alpha$ in $(\alpha,x)$-robustness, can grow larger slowly as the game grows large and the Price of Anarchy will still tend to one.
@@
Define the spread of the game to be $T\coloneqq\frac{W_{1}}{max_{i,k}v_{k}^{i}}$. Suppose we run SQUAP with an aggregation mechanism that is alternative-independent and $(\alpha,x)$-robust. Then in the two-alternative setting, there exists a choice of $c$ such that
$\text{pPoA}(\mathbf{v},\mathcal{D})\geq 1-\frac{2\alpha}{T}-\left(\frac{4}{T}\right)^{2/5}.$
The general strategy of proving Theorem 2 mirrors our QTM analysis, where we consider the case of a large welfare gap $W_{1}-W_{2}$. We first prove:
@@
If the aggregation mechanism is alternative-independent and $(\alpha,x)$-robust, and $W_{1}-W_{2}\geq 2\alpha\max_{i,k}v_{k}^{i}$, then in any submechanism-perfect pure equilibrium of the two alternative setting,
@@
Define $\hat{W}_{k}=V_{k}+\hat{B}_{k}$, and recall $W_{k}=V_{k}+B^{*}_{k}$. By Lemma 1, we have
$\hat{W}_{1}-\hat{W}_{2}$ $=W_{1}-W_{2}+\hat{B}_{1}-B^{*}_{1}+\hat{B}_{2}-B^{*}_{2}$
@@
In the QTM stage, the synthetic players’ aggregate values are $\mathbf{\hat{B}}$. As mentioned before, under alternative independence forecasters will not manipulate their votes to increase utility in the aggregation stage. Thus, agents in equilibrium will vote strategically in the QTM stage exactly as in Section 4. It follows by Corollary 2 that
@@
$\geq 1-\left(\frac{8c}{Y}\right)^{2/3}$
@@
Equipped with Lemma 1, we can also give our Price of Anarchy result for $m$ alternatives. This result is phrased as a reduction from any Price of Anarchy bound for the QTM under the same scenario but without predictions. The proof is similar to the above and appears in Appendix B.
Theorem 3.
Define the spread of the game to be $T\coloneqq\frac{W_{1}}{max_{i,k}v_{k}^{i}}$. Suppose we run SQUAP with an aggregation mechanism that is alternative-independent and $(\alpha,x)$-robust, and suppose the QTM with $m$ alternatives has a Price of Anarchy guarantee $\text{pPoA}{(\mathbf{v})}\geq C(\mathbf{v})$. Then the Price of Anarchy of SQUAP satisfies
$\text{pPoA}(\mathbf{v},\mathcal{D})\geq\frac{1}{1+\alpha}\left(C(\mathbf{v})-\frac{\alpha}{T}\right).$
We now have a Price of Anarchy bound of Theorem 2 for any information-aggregation mechanism that is alternative-independent and $(\alpha,x)$-robust. Next, we show that two such information aggregation mechanisms do actually exist.
5.3 Wagering Mechanisms
We consider using wagering mechanisms as the information-aggregation mechanism in stage 1 of SQUAP. Wagering mechanisms are a common tool for eliciting and aggregating predictions (Lambert et al., 2008, 2015b). For simplicity, we will suppose that agents predict $\mathbf{B}$ directly and the mechanism is later able to exactly observe $B_{k}$, the true external welfare of the alternative $k$ that was selected. When observing $B_{k}$ is difficult, one could consider wagers that are more complex and predict other relevant variables. In this case, the mechanism could compute an estimate of $\mathbf{B}$ from the mechanism’s forecasts and use the estimate in SQUAP.
We consider a natural extension that does not appear to have been studied: decision wagering mechanisms. Instead of agents wagering on a single future event, we will have them wager on the future event conditional on alternative $k$ being selected, for each $k=1,\dots,m$. Once $k$ is selected and the event is observed, we cancel all other wagers and assign payments based on the predictions for $k$.
Specifically, we adapt the Brier betting mechanism (Lambert et al., 2015a) to our setting, defining the importance-weighted quadratic decision wagering mechanism. The Brier betting mechanism makes use of the quadratic score $s$ with parameter $\beta>0$:
$s(\hat{b},b)=-\frac{1}{\beta}(\hat{b}-b)^{2}.$ (5)
$s$ is one example of a proper scoring rule for the mean (Savage, 1971; Gneiting and Raftery, 2007): the expected value of $s(\hat{b},b)$, over draws $b$ from some distribution $\mathcal{D}$, is maximized by choosing $\hat{b}=\operatorname{\mathbb{E}}[b]$. Following the prediction market literature, we refer to $\beta$ as the liquidity parameter. $\beta$ determines how sensitive the payoffs are to small changes in a prediction. If $\beta$ is small, then the total payoffs available are very large, but a significant payment may be required as well.
Formally, for each alternative $k$, forecaster $i$ submits an estimate $\hat{b}^{i}_{k}$ for the expected external welfare conditioned on that alternative. Let $\hat{\mathbf{b}}$ consist of all the players’ predictions for all the alternatives. Once an alternative $k\sim\mathbf{p}$ is chosen and the outcome $b_{k}$ is observed, each player $i$ receives the payoff
$\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k})=\frac{1}{p_{k}}\left[s(\hat{b}^{i}_{k},b_{k})-\frac{1}{n}\sum_{j=1}^{n}s(\hat{b}^{j}_{k},b_{k})\right].$
In other words, forecaster $i$’s payoff consists of her score minus the average score of all participants. Since $\sum_{i}\pi_{i}(\mathbf{\hat{b}},\mathbf{p},k,b_{k})=0$, we obtain:
Observation 1.
Wagers and agent loss.
Above, we have considered the special case of a wagering mechanism where all wagers are equal to one. In general, the wagering mechanism weights each participant’s score $s(\hat{b}^{i}_{k},b_{k})$ by $\frac{y_{i}^{k}}{Y_{k}}$, where $y^{i}_{k}$ is some nonnegative wager on alternative $k$ and $Y_{k}=\sum_{i}y^{i}_{k}$. The wagering mechanism is generally applied in cases where scores $s()$ are bounded in $[0,1]$, so that an agent can be guaranteed to win or lose no more than her wager. In our case, such guarantees can be achieved by adjusting $\beta$, the liquidity parameter of $s$, if an upper bound on $B_{k}$ is known. Wagering mechanisms can also be subsidized, e.g. by additionally giving each agent a fraction of their score, or by making all payouts nonnegative with a shift. We could use a similar analysis to the prediction market setting to use the QTM revenue to subsidize the wagering mechanism (see Corollary 5.)
Strategic behavior and aggregation.
For analysis of this mechanism, unlike with prediction markets in Section 5.4, we will need to adopt the standard wagering mechanism model of immutable beliefs, where each agent $i$ has fixed beliefs $\mathbf{b}^{i}$ about the expectation of $\mathbf{B}$, regardless of others’ beliefs. (In a Bayesian model, we essentially have a “no-trade” situation in which the zero-sum payoffs discourage participation.) The wagering mechanism in isolation is strategyproof: agents maximize expected net payoff by reporting their true believed estimates. This follows immediately from the properness of the scoring rule $s$ and the fact that in $\pi_{i}$, the agent can only affect their own score. For analysis, we suppose that the wagering mechanism outputs $\hat{B}_{1},\dots,\hat{B}_{m}$ where $\hat{B}_{k}=\frac{1}{N}\sum_{i=1}^{N}\hat{b}_{k}^{i}$. We assume that, if all agents are truthful, then the output $\mathbf{B^{*}}$ is the true expectation of the external welfare impacts. Our analysis could be adapted to other aggregation methods than taking the average.
Now consider agent $i$. The only terms in her expected utility that $i$ controls are
$\frac{1}{p_{k}}\left(s(\hat{b}_{k}^{i},b_{k})-\frac{1}{n}s(\hat{b}_{k}^{i},b_{k})\right)=\left(\frac{n-1}{n}\right)\frac{1}{p_{k}}s(\hat{b}_{k}^{i},b_{k})$
for each $k$. If the outcome $b_{k}$ is drawn from a distribution $\mathcal{D}_{k}$, then the expected score over both $k\sim\mathbf{p}$ and $b_{k}\sim\mathcal{D}_{k}$ is
$\sum_{k}p_{k}\operatorname{\mathbb{E}}_{\mathcal{D}_{k}}\left[\left(\frac{n-1}{n}\right)\frac{1}{p_{k}}s(\hat{b}_{k}^{i},b_{k})\right]$ $=-\frac{n-1}{\beta n}\sum_{k}\left((\hat{b}_{k}^{i}-b_{k}^{i})^{2}+\text{Var}(D_{k})\right).$
The variance term is independent of any agents’ actions, and moreover cancels out in an agent’s payoff when we take the difference over different strategies. Therefore, it is without loss of generality to define the expected score of a prediction $\mathbf{\hat{b}}$ to be
$S(\mathbf{\hat{b}}^{i};\mathbf{b}^{i})$ $=-\frac{n-1}{\beta n}\sum_{k}(\hat{b}_{k}^{i}-b_{k}^{i})^{2}.$ (6)
Because (6) does not depend on the selected outcome $k$, we obtain:
Observation 2.
The importance-weighted quadratic decision wagering mechanism with the importance-weighted quadratic score is alternative-independent.
Price of Anarchy.
We now have the tools to analyze SQUAP with the importance-weighted quadratic decision wagering mechanism.
Lemma 3.
The importance-weighted quadratic decision wagering mechanism with parameter $\beta=\frac{1}{2}\epsilon x$ is $(\epsilon^{1/2},x)$-robust.
Proof.
By Equation 6, the loss from misreporting $\mathbf{\hat{b}}^{i}$ is
$S(\mathbf{\hat{b}}^{i};\mathbf{b}^{i})-S(\mathbf{b}^{i};\mathbf{b}^{i})$ $=-\frac{n-1}{\beta n}\sum_{k}(\hat{b}_{k}^{i}-b_{k}^{i})^{2}$
$\leq-\frac{1}{2\beta}\sum_{k}(\hat{b}_{k}^{i}-b_{k}^{i})^{2},$
where the last line follows since $n\geq 2$.
By the definition of $x$-best-responding in the wagering mechanism, $x>\frac{1}{2\beta}\sum_{k}(\hat{b}_{k}^{i}-b_{k}^{i})^{2}$. So for all $k$, $(\hat{b}_{k}-b_{k}^{i})^{2}\leq 2\beta x=\epsilon x^{2}$, or $|\hat{b}_{k}^{i}-b_{k}^{i}|\leq\epsilon^{1/2}x$. Now, this holds for all participants, and the output of the wagering mechanism is $\hat{B}_{k}=\frac{1}{n}\sum_{i=1}^{n}\hat{b}^{i}_{k}$, so by definition the mechanism is $(\epsilon^{1/2},x)$-robust. ∎
Corollary 3.
Let $T=\frac{W_{1}}{\max_{i,k}v_{k}^{i}}$ be the spread. In the two alternative setting, SQUAP with the importance-weighted quadratic decision mechanism with wagering parameter $\beta=\frac{1}{2}\epsilon\max_{i,k}v_{k}^{i}$ and QTM parameter $c=\tfrac{1}{2}\max_{i,k}v_{k}^{i}$ satisfies
$\text{pPoA}(\mathbf{v},\mathcal{D})\geq 1-\frac{2\epsilon^{1/2}}{T}-\left(\frac{4}{T}\right)^{2/5}.$
We observe that, even with constant liquidity $\epsilon=\Theta(1)$, the Price of Anarchy tends to one with the spread. When $m>2$ and the QTM with $m$ alternatives has a Price of Anarchy guarantee $\geq C(\mathbf{v})$, we could also apply Theorem 3 to recover a Price of Anarchy guarantee $\frac{1}{1+\epsilon^{1/2}}\left(C(\mathbf{v})-\frac{\epsilon^{1/2}}{T}\right)$.
@@ 5.4 Prediction Markets
Next we consider using prediction markets as the information-aggregation mechanism in stage $1$ of SQUAP. Prediction markets may be considered more complex than wagering mechanisms since they are not one-shot. However, while the designer must combine the resulting predictions from a wagering mechanism herself (often by averaging), prediction markets automatically aggregate information.
Prediction markets are financial markets designed specifically for aggregating the predictions of agents by rewarding accuracy once an outcome is observed. We briefly define scoring-rule based prediction and decision markets below, referring the reader to references in Section 1.2 for more background.
@@
We define a scoring-rule prediction market for a real-valued random variable, following e.g. Abernethy et al. (2013). The market defines an initial estimate $\hat{b}^{0}$. Agents one at a time arrive in sequence and provide estimates $\hat{b}^{1},\dots,\hat{b}^{N}$. Later, when the outcome $b$ of the random variable is observed, the provider of each estimate $\hat{b}^{t}$ receives a net payoff $s(\hat{b}^{t},b)-s(\hat{b}^{t-1},b)$, where $s$ is some proper scoring rule as in the wagering mechanism setting. In particular, we will again focus on the quadratic scoring rule as defined in Equation 5.
Observe that each participant’s net payoff can be positive or negative, and the total payment of the mechanism telescopes to $s(\hat{b}^{N},b)-s(\hat{b}^{0},b)$. Following the prediction market literature and our section on wagering mechanisms, $\beta$ represents the liquidity parameter. We have defined the market so that agents arrive only once each, but our results are robust to agents participating multiple times or at strategically chosen times, subject to an assumption discussed next.
Solution concept.
Since the prediction market framework is not a one-shot game, but instead an extensive-form Bayesian game, we must modify our solution concept from subgame perfect equilibrium. We say the combined mechanism is in sub-mechanism perfect equilibrium if it is in Nash equilibrium and, in addition, the strategies played at each subgame consisting of the Synthetic Players QTM are in pure-strategy Nash equilibrium. However, we will need further assumptions on the strategies played in the game as well.
Strategic behavior in prediction markets is highly complex and not fully understood (Chen et al., 2010; Ostrovsky, 2012; Gao et al., 2013; Chen and Waggoner, 2016). However, one robust result of all of the above works is that, in the long run, agents do not stay misinformed. They may be under-informed, but indeed all information is often aggregated in equilibrium (with an appropriate solution concept such as Weak Perfect Bayesian equilibrium). In particular, it is known that in strategic equilibrium of prediction markets, information is always aggregated under a condition called separable securities (Ostrovsky, 2012). In other words, when information aggregation fails in prediction markets, it is not for strategic reasons, but only because of the expressiveness of the prediction language. Therefore, we believe that efficient markets is a mild assumption on strategic behavior of the agents in the combined mechanism, and it can be achieved by introducing additional tradeable securities.
Definition 3.
We say the efficient markets assumption holds on the decision market if, in any equilibrium, in any subgame (including off the equilibrium path), the last participant in the decision market knows and believes $\mathbf{B^{*}}$, the true expected external welfare impacts of the alternatives conditioned on all available information.
In other words, we assume that a manipulator is not able to cause information to be hidden or the other participants to be misled. However, a manipulator can still change the final predictions by waiting until the final stage to trade.
The impact of the efficient markets assumption is that a manipulator is best off if they are the one predicting last. She waits for the markets to converge to the prediction $\mathbf{\hat{B}}=\mathbf{B^{*}}$, then manipulates $\mathbf{\hat{B}}$ arbitrarily. We will show that such a manipulation will not be large enough to damage the welfare guarantee of SQUAP.
@@
In a decision market (e.g. Chen et al. (2011)), we simultaneously run $m$ prediction markets, one for each of $m$ alternatives. Here, we model this as each participant $t=1,\dots$ providing an estimate $\mathbf{\hat{b}}^{t}\in\operatorname{\mathbb{R}}^{m}$. Then, we make a decision $k$ based on the outcome of the markets, and later observe $b_{k}$. We cancel all trades in the other $m-1$ markets, i.e. we only assign payoffs based on predictions $\hat{b}_{k}^{t}$ made in the $k$th prediction market.
As we saw with decision wagering mechanisms, decision markets can lead to incentives to misreport and manipulate the decision. We thus use the same importance-weighted scoring rule proposed by Chen et al. (2011):
$S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k})=\frac{1}{p_{k}}s(\hat{b}^{t}_{k},b_{k}).$
Then if the outcome $b_{k}$ is drawn from a distribution $D_{k}$, the expected score over both $k\sim\mathbf{p}$ and $b_{k}\sim D_{k}$ is
$\sum_{k}p_{k}\operatorname{\mathbb{E}}_{\mathcal{D}_{k}}S(\mathbf{\hat{b}}^{t},\mathbf{p},k,b_{k})$ $=\sum_{k}\operatorname{\mathbb{E}}_{\mathcal{D}_{k}}(\hat{b}_{k}^{t}-b_{k})^{2}$
$=\frac{-1}{\beta}\sum_{k}\left((\hat{b}_{k}^{t}-b^{*}_{k})^{2}+\text{Var}(\mathcal{D}_{k})\right),$
where $b_{k}^{*}=\operatorname{\mathbb{E}}_{D_{k}}[b_{k}]$. For the same reasons as before, we can ignore the variance term since it is independent of agent actions and cancels out in the market scoring rule. Therefore, under the importance-weighted scoring rule, it is without loss of generality to define the expected score of a prediction $\mathbf{\hat{b}}$, given the vector of means $\mathbf{b^{*}}=b_{1}^{*},\dots,b_{m}^{*}$, to be
$S(\mathbf{\hat{b}};\mathbf{b^{*}})$ $=\frac{-1}{\beta}\sum_{k}(\hat{b}_{k}-b_{k}^{*})^{2}.$ (7)
We obtain:
Observation 3.